### v2

# HD2-SSC: High-Dimension High-Density Semantic Scene Completion for Autonomous Driving

Abstract
Camera-based 3D semantic scene completion (SSC) plays a crucial role in autonomous driving, enabling voxelized 3D scene understanding for effective scene perception and decision-making. Existing SSC methods have shown efficacy in improving 3D scene representations, but suffer from the inherent input-output dimension gap and annotation-reality density gap, where the 2D planar view from input images with sparse annotated labels leads to inferior prediction of real-world dense occupancy with a 3D stereoscopic view. In light of this, we propose the corresponding High-Dimension High-Density Semantic Scene Completion (HD2-SSC) framework with expanded pixel semantics and refined voxel occupancies. To bridge the dimension gap, a High-dimension Semantic Decoupling module is designed to expand 2D image features along a pseudo third dimension, decoupling coarse pixel semantics from occlusions, and then identify focal regions with fine semantics to enrich image features. To mitigate the density gap, a High-density Occupancy Refinement module is devised with a “detect-and-refine” architecture to leverage contextual geometric and semantic structures for enhanced semantic density with the completion of missing voxels and correction of erroneous ones. Extensive experiments and analyses on the SemanticKITTI and SSCBench-KITTI-360 datasets validate the effectiveness of our HD2-SSC framework.
Codehttps://github.com/PKU-ICST-MIPL/HD2-AAAI2026
Figure 1: Illustrations of (a) Dimension Gap, highlighting the disparity between input coarse pixel semantics with occlusion and output distinct fine voxel semantics, (b) Density Gap, depicting the difference between annotated sparse labels and ground-truth dense occupancies.
Introduction
Accurate 3D geometry perception of the surrounding environment is essential for autonomous driving systems, enabling downstream tasks such as planning, navigation, and interaction with dynamic environments. 3D Semantic Scene Completion (SSC) (roldao20223d) aims to infer both the occupancy and semantics of the 3D space, providing a voxelized understanding of the scene. Considering the complexity of outdoor environments, LiDAR-based methods (roldao2020lmscnet; cheng2021s3cnet; xia2023scpnet; li2023lode; lee2024semcity; wang2025l2cocc) have been widely explored due to their ability to directly capture 3D structural information from input point clouds. However, they face the inherent challenges of high sensor cost and limited scalability for LiDAR data, restricting their applications to large-scale deployment in real-world scenarios.
Recently, camera-based semantic scene completion has garnered increasing attention as a cost-effective and scalable alternative utilizing 2D images as input. The pioneering work MonoScene (cao2022monoscene) lifts 2D image features into dense 3D volumes with depth projection. Building on this, bird’s-eye-view (BEV) (huang2021bevdet; li2023bevstereo) and tri-perspective view (TVP) (huang2023tri) representations are designed for improved computation efficiency. With these advances, transformer-based methods (li2022bevformer; wei2023surroundocc) utilized voxel queries to aggregate image features into 3D scenes. VoxFormer (li2023voxformer) introduces a two-stage framework, aggregating 3D features of visible areas and then diffusing to the whole scene. Efforts have been made for improved performance through neighboring object structures and view spatial differences (xue2024bi), context-dependent queries (yu2024context), and dense-sparse-dense design for hybrid guidance (mei2024camera).
Despite significant progress, existing methods typically focus on refining voxel-based scene representations, but treat pixel features and voxel semantics indiscriminately during view transformation and occupancy prediction, facing two key challenges. (1) Dimension Gap: As illustrated in Figure 1 (a), input images are captured from a 2D planar view, leading to coarse pixel features confused by multiple object semantics with occlusions. In contrast, SSC requires fine-grained voxel semantics for distinct objects in a 3D stereoscopic view, necessitating the expansion and decoupling of coarse pixel features from occlusion. (2) Density Gap: As shown in Figure 1 (b), manual annotations derived from LiDAR sensors yield inherently sparse labels with interspace due to the finite resolution of collected point clouds. On the other hand, real-world scenes exhibit dense, consistent occupancy with rich contextual details, demanding comprehensive detection and meticulous refinement of voxel occupancy to enrich semantic density.
To address these challenges, we introduce the High-Dimension High-Density Semantic Scene Completion (HD2-SSC) framework. Specifically, the High-dimension Semantic Decoupling (HSD) module first expands and decouples coarse pixel semantics to alleviate the dimension gap. The process begins with expanding 2D image features into voxelized features through the Pseudo Voxelization block, with an orthogonal loss promoting distinct expanded semantics. The Semantic Aggregation block then extracts global semantics with pixel queries and integrates high-dimension voxelized features using semantic clustering with decoupling loss. Then, to mitigate the density gap, the High-density Occupancy Refinement (HOR) module operates in a “detect-and-refine” manner to enhance semantic density with contextual geometric and semantic structures. In the detection phase, binary predictions are generated for a comprehensive collection of occupied voxels, focusing on the identification of geometric critical voxels. In the subsequent refinement phase, we conduct class-wise prediction to identify semantic critical voxels, with respect to the prediction confidence score. Then the overall distributions of geometric and semantic critical voxels are aligned for consistent contextual structures, enhancing semantic density with completion of missing voxels and correction of erroneous ones. Extensive experiments and analyses on SemanticKITTI and SSCBench-KITTI-360 demonstrate the superiority of our HD2-SSC over SOTA methods.
The main contributions are summarized as follows:
  • $\bullet$
    We propose HD2-SSC to address the challenges of dimension and density gap in camera-based semantic scene completion, improving SSC performance with completion of missing voxels and correction of erroneous ones.
  • $\bullet$
    HSD bridges the dimension gap through expanding and decoupling coarse pixel semantics with orthogonal loss, then aggregating high-dimension voxelized semantics upon semantic clustering with decoupling loss.
  • $\bullet$
    HOR addresses the density gap in a “detect-and-refine” paradigm to identify geometric and semantic critical voxels, aligning the overall critical distributions for consistent contextual details and improved semantic density.
Figure 2: The overall architecture of our HD2-SSC. The High-dimension Semantic Decoupling (HSD) module expands and decouples coarse pixel semantics with orthogonal loss, then aggregates high-dimension voxelized semantics via semantic clustering with decoupling loss. The High-density Occupancy Refinement (HOR) module adopts a “detect-and-refine” architecture to identify geometric and semantic critical voxels, whose overall distributions are aligned for consistent contextual details.
Related Work
LiDAR-based SSC Methods
Due to the complexity of outdoor scenes and rich 3D structural information embedded in point clouds, LiDAR-based methods (liong2020amvnet; cheng20212; zhang2023attention; ye2023lidarmultinet) have long been the predominant solutions for SSC. UDNet (zou2021up) pioneered the use of a 3D U-Net architecture to generate scene predictions directly from LiDAR data. SGCNet (zhang2018efficient) enhanced efficiency through spatial group convolutions for sparse 3D processing. Further advancements have focused on refining 3D scene representations, such as multi-view fusion for completing sparse scenes (cheng2021s3cnet) and local implicit functions for continuous scene modeling (rist2021semantic). SCPNet (xia2023scpnet) transferred multi-frame information to single-frame models for lightweight architectures. SSA-SC (yangsemantic) and SSC-RS (mei2023ssc) adopted multi-branch designs to hierarchically fuse semantic and geometric features. Despite the advancements, LiDAR-based methods remain computationally expensive due to large volume of LiDAR points, limiting their scalability to real-world, large-scale applications.
Camera-based SSC Methods
In recent years, camera-based SSC methods (miao2023occdepth; yao2023ndc; tong2023scene; wang2024h2gformer; wang2024not; yang2025sphere; yang2025gala) have gained increasing attention due to their cost-effectiveness and ease of deployment. One of the first fully visual SSC frameworks, MonoScene (cao2022monoscene), pioneers a voxel-based approach by sampling image features along lines of sight. Another line of research leverages Bird’s-Eye-View (BEV) representations to aggregate multi-view image features for 3D scene modeling. LSS (philion2020lift) introduces a frustum-based lifting mechanism, which is further enhanced by BEVDet (huang2021bevdet), aligning multi-frame features for improved scene comprehension. BEVDepth (li2023bevdepth) and BEVStereo (li2023bevstereo) incorporate camera-aware depth estimation and dynamic temporal stereo information for more precise depth learning. Beyond BEV, TPVFormer (huang2023tri) proposes a Tri-Perspective View (TPV) representation, decomposing voxel grids onto three orthogonal planes for efficient scene encoding. Transformer-based methods utilize BEV queries (jiang2023polarformer) or voxel queries (ma2024cotr). Voxformer (li2023voxformer) adopts a two-stage framework with class-agnostic query proposals. Symphonize (jiang2024symphonize) introduces a scene-from-instance paradigm to enhance feature interactions. HASSC (wang2024not) exploits the local and global hardness of each voxel, together with a self-distillation strategy for stable and consistent training process. SGN (mei2024camera) presents a one-stage SSC framework with a dense-sparse-dense strategy, exploiting hybrid guidance to improve segmentation boundaries for more precise 3D semantic scene completion.
In general, existing camera-based SSC methods mainly focus on devising delicate architectures to refine 3D features, but overlook two critical challenges: the input-output dimension gap and annotation-reality density gap. In contrast, our HD2-SSC approach decouples coarse pixel semantics to bridge the dimension gap for aligned view transformation, and aligns critical voxel distributions for consistent contextual details to mitigate the density gap.
Approach
Figure 2 illustrates our High-Dimension High-Density 3D Semantic Scene Completion (HD2-SSC) framework, consisting of four main modules: (1) an Image Encoder extracting 2D features, (2) a High-dimension Semantic Decoupling (HSD) module that addresses the dimension gap by decoupling coarse pixel semantics with orthogonal loss and aggregating high-dimension voxelized semantics upon semantic clustering, (3) a View Transform module projecting 2D pixel features to 3D voxel features, and (4) a High-density Occupancy Refinement (HOR) module that mitigates the density gap through aligning geometric and semantic critical voxel distributions in a “detect-and-refine” manner.
Preliminary
Problem Setup.
Given input stereo images $I_{l}^{\rm rgb},I_{r}^{\rm rgb}$, the objective of SSC is to predict the geometry and semantics of 3D scenes in front. The output is represented as a voxel grid $Y\in\mathbb{R}^{H\times W\times Z}$, where $H,W,Z$ correspond to the grid’s length, width, and height, respectively. Each voxel is categorized as either empty denoted by $c_{0}$ or occupied by one of the semantic classes in $c\in\{c_{1},\cdots,c_{N}\}$, where $N$ represents the number of semantic classes.
Image Encoder.
Consistent with previous methods (mei2024camera), we employ the ResNet-50 (he2016deep) network with FPN (lin2017feature) to extract 2D features $F^{\rm 2D}\in\mathbb{R}^{N_{t}\times C\times H\times W}$ from input RGB images, where $N_{t}$ is the image number of temporal inputs, $C$ is the feature channel and $(H,W)$ denotes the image resolution.
View Transformer.
Following (mei2024camera), we construct 3D features by sampling 2D features via 3D-2D projection mapping with camera parameters. Let $x\in\mathbb{R}^{H\times W\times Z\times 3}$ denote the centroid of $H\times W\times Z$ voxels in world coordinates. We establish the projection mapping $\pi(x)$ from pixel $(u,v)$ using the camera intrinsic parameter $K$ and extrinsic parameter $T=[R,t]$:
T $=R\cdot p+t$ (1)
$z_{c}\circ[u,v,1]^{T}$ $=K\cdot[h_{c},w_{c},z_{c}]^{T}$
where $\circ$ denotes element-wise product.
High-dimension Semantic Decoupling
The High-dimension Semantic Decoupling (HSD) module consists of two blocks for addressing the dimension gap. The Pseudo Voxelization block expands and decouples coarse pixel semantics, and the Semantic Aggregation block aggregates the high-dimension voxelized semantics.
Pseudo Voxelization.
Given image feature $F_{\rm cam}$ extracted from the image encoder, a Dim Expansion (DE) layer is employed to lift it along a pseudo “semantic dimension” into pseudo voxelized features. This step is essential for addressing the dimension gap, generating multiple candidates for decoupling occluded objects within coarse pixel semantics:
$F_{\rm pseudo}={\rm DE}(F_{\rm cam},D_{\rm exp})$ (2)
where $F_{\rm pseudo}=\{F_{\rm pseudo}^{i}\in\mathbb{R}^{N_{\rm cam}\times C_{\rm 2D}\times H_{\rm 2D}\times W_{\rm 2D}}\}_{i=1}^{D_{\rm exp}}$ represents pseudo voxelized features, and $D_{\rm exp}$ is the channel of expanded dimension, ${\rm DE}(\cdot)$ represents the dim expansion operation with 2D convolution layers. However, the pseudo voxelized features may still exhibit similar semantics at the same 2D coordinates $(h_{\rm 2D},w_{\rm 2D})$, which limits their ability to generalize across different occluded objects. Therefore, we employ an orthogonal loss to further promote the distinct voxelized semantics:
$\mathcal{L}_{\rm orth}=\lambda\left|W_{\rm DE}W_{\rm DE}^{\rm T}-I\right|$ (3)
where $W_{\rm DE}$ represents the weight matrix of the dim expansion projection, $I$ indicates the identity matrix, and $\lambda$ is the regularization parameter for the orthogonal loss.
Semantic Aggregation.
We first introduce pixel queries $Q_{\rm pixel}\in\mathbb{R}^{N_{\rm query}\times C_{\rm 2D}}$ to collect global semantics from the pseudo voxelized features through cross attention mechanism. Then, to further assign different pseudo voxelized feature slices $F_{\rm pseudo}^{i}$ with discrete semantics, we perform semantic clustering (DPC-kNN) on the collected global semantics in $Q_{\rm pixel}$:
[…]
where $C$ represents the $D_{\rm exp}$ semantic clusters corresponding to the expanded dimension of the pseudo voxelized features, and $c_{i}$ is the centroid of the $i$-th cluster $C_{i}$. Additionally, a decoupling loss is adopted to enhance the distinction among different semantic clusters:
$\mathcal{L}_{\rm decouple}=\sum\limits_{i\neq j}\dfrac{C_{i}\cdot C_{j}}{|C_{i}|\cdot|C_{j}|}$ (5)
Then, discriminative regions of the pseudo voxelized features are located with respect to the similarity against semantic clusters, thereby aggregating high-dimension voxelized semantics, as illustrated in Figure 3:
$F_{\rm agg}=\sum\limits_{i=1}^{D_{\rm exp}}F_{\rm pseudo}^{i}\cdot\mathop{\max}\limits_{C_{j}}{\rm sim}(F_{\rm pseudo}^{i},C_{j})$ (6)
Figure 3: Illustration of aggregating the high-dimension voxelized semantics concerning the semantic clusters with decoupling loss.
High-density Occupancy Refinement
The High-density Occupancy Refinement (HOR) module is designed in a “detect-and-refine” architecture to identify geometric and semantic critical voxels, respectively. Then the critical voxels are aligned to ensure contextual geometric and semantic consistency.
Method SC SSC
$\blacksquare$ car(3.92%)
$\blacksquare$ bicycle(0.03%)
$\blacksquare$ motorcycle(0.03%)
$\blacksquare$ truck(0.16%)
$\blacksquare$ other-veh.(0.20%)
$\blacksquare$ person(0.07%)
$\blacksquare$ bicyclist(0.07%)
$\blacksquare$ motorcyclist(0.05%)
$\blacksquare$ road(15.30%)
$\blacksquare$ parking(1.12%)
$\blacksquare$ sidewalk(11.13%)
$\blacksquare$ other-grnd.(0.56%)
$\blacksquare$ building(14.10%)
$\blacksquare$ fence(3.90%)
$\blacksquare$ vegetation(39.3%)
$\blacksquare$ trunk(0.51%)
$\blacksquare$ terrain(9.17%)
$\blacksquare$ pole(0.29%)
$\blacksquare$ traf.-sign(0.08%)
IoU mIoU
Mono-Input Methods
MonoScene (cao2022monoscene) 36.86 11.08 23.26 0.61 0.45 6.98 1.48 1.86 1.20 0.00 56.52 14.27 26.72 0.46 14.09 5.84 17.89 2.81 29.64 4.14 2.25
TPVFormer (huang2023tri) 35.61 11.36 23.81 0.36 0.05 8.08 4.35 0.51 0.89 0.00 56.50 20.60 25.87 0.85 13.88 5.94 16.92 2.26 30.38 3.14 1.52
OccFormer (zhang2023occformer) 36.50 13.46 25.09 0.81 1.19 25.53 8.52 2.78 2.82 0.00 58.85 19.61 26.88 0.31 14.40 5.61 19.63 3.93 32.62 4.26 2.86
IAMSSC (xiao2024instance) 44.29 12.45 26.26 0.60 0.15 8.74 5.06 1.32 3.46 0.01 54.55 16.02 25.85 0.70 17.38 6.86 24.63 4.95 30.13 6.35 3.56
Temporal-Input Methods
VoxFormer (li2023voxformer) 44.15 13.35 26.54 1.28 0.56 7.26 7.81 1.93 1.97 0.00 53.57 19.69 26.52 0.42 19.54 7.31 26.10 6.10 33.06 9.15 4.94
DepthSSC (yao2023depthssc) 45.84 13.28 25.94 0.35 1.16 6.02 7.50 2.58 6.32 0.00 55.38 18.76 27.04 0.92 19.23 8.46 26.37 4.52 30.19 7.42 4.09
Symphonies (jiang2024symphonize) 41.92 14.89 28.68 2.54 2.82 20.44 13.89 3.52 2.24 0.00 56.37 15.28 27.58 0.95 21.64 8.40 25.72 6.60 30.87 9.57 5.76
HASSC (wang2024not) 44.58 14.74 27.33 1.07 1.14 17.06 8.83 2.25 4.09 0.00 57.23 19.89 29.08 1.26 20.19 7.95 27.01 7.71 33.95 9.20 4.81
H2GFormer (wang2024h2gformer) 44.69 14.29 28.21 0.95 0.91 6.80 9.32 1.15 0.10 0.00 57.00 21.74 29.37 0.34 20.51 7.98 27.44 7.80 36.26 9.88 5.81
CGFormer (yu2024context) 45.99 16.87 34.32 4.61 2.71 19.44 7.67 2.38 4.08 0.00 65.51 20.82 32.31 0.16 23.52 9.20 26.93 8.83 39.54 10.67 7.84
SGN (mei2024camera) 46.21 15.32 33.31 0.61 0.46 6.03 9.84 0.47 0.10 0.00 59.10 19.05 29.41 0.33 25.17 9.96 28.93 9.58 38.12 13.25 7.32
HD2-SSC (Ours) 47.59 17.44 32.68 1.72 0.54 24.76 10.73 1.61 1.73 0.00 60.80 23.72 33.12 0.11 28.55 8.83 30.90 11.08 38.78 13.41 8.34
Table 1: Camera-based 3D semantic scene completion results on the SemanticKITTI (behley2019semantickitti) validation set. denotes the methods employing EfficientNet-B7 (tan2019efficientnet) as image backbone, and represents the methods utilizing ResNet-50 (he2016deep) as image backbone. Best results are highlights in bold, and second-best scores are underlined.
Detection Phase.
During the detection phase, the voxel queries $Q_{\rm voxel}\in\mathbb{R}^{N_{\rm query}\times C_{\rm 3D}}$ and voxel feature $F_{\rm voxel}\in\mathbb{R}^{C_{\rm 3D}\times H_{\rm 3D}\times W_{\rm 3D}\times Z_{\rm 3D}}$ are passed through a binary classification head for a comprehensive detection of occupied voxels, together with a heuristic separation between foreground and background voxels:
$M_{\rm o-f},M_{\rm f-b}=H_{\rm bc}(F_{\rm voxel},Q_{\rm voxel})$ (7)
where $H_{\rm bc}$ denotes the binary classification head, $M_{\rm o-f},M_{\rm f-b}$ represent the score maps distinguishing occupied voxels from free ones, foreground voxels from background ones, respectively. The generated score maps are added to represent the geometric density score of each voxel, indicating its general structural significance. Then we select geometric critical voxels with the highest geometric density scores:
$V_{\rm geo}={\rm top}_{k}(M_{\rm o-f}+M_{\rm f-b})$ (8)
where $k$ is the number of selected geometric critical voxels.
In the detection phase, we focus on the general classification between occupied and free, foreground and background voxels, and select geometric critical voxels with discriminative structural contextual information.
Refinement Phase.
During the refinement phase, we first conduct class-wise prediction with the voxel queries and voxel features to generate the initial semantic scene completion results:
$Y_{\rm init}=H_{\rm cc}(F_{\rm voxel},Q_{\rm voxel})$ (9)
where $H_{\rm cc}$ denotes the class-wise classification head, $Y_{\rm init}$ represents the initial semantic scene completion results. Furthermore, we select semantic critical voxels with $Y_{\rm init}$ indicating the classification confidence and contextual semantic significance:
$V_{\rm sem}={\rm top}_{k}\left(\max(Y_{\rm init})\right)$ (10)
where $k$ is the number of selected semantic critical voxels, aligned with geometric critical voxels.
Voxel Alignment.
After selecting geometric and semantic critical voxels with discriminative geometric and semantic contextual information, respectively, we align the overall critical voxel distributions to promote consistent contextual semantic and geometry for further refinement:
$\mathcal{L}_{\rm critical}=\mathcal{L}_{\rm KL}(V_{\rm geo}||V_{\rm sem})+\mathcal{L}_{\rm KL}(V_{\rm sem}||V_{\rm geo})$ (11)
where $\mathcal{L}_{\rm KL}(\cdot||\cdot)$ is the Kullback-Leibler divergence loss measuring the difference between two distributions. With the aligned critical voxels serving as soft constraint on the complementary effect of contextual geometry and semantics, we utilize a refined multi-layer perceptron ($\rm MLP_{refine}$}) to enhance the initial SSC predictions:
$Y_{\rm refine}=Y_{\rm init}+{\rm MLP_{refine}}([V_{\rm geo},V_{\rm sem}])$ (12)
where $Y_{\rm refine}$ represents the final refined SSC prediction, $\cdot$$\cdot$] is the concatenation operation.
Method SC SSC
$\blacksquare$ car(2.85%)
$\blacksquare$ bicycle(0.01%)
$\blacksquare$ motorcycle(0.01%)
$\blacksquare$ truck(0.16%)
$\blacksquare$ other-veh.(5.75%)
$\blacksquare$ person(0.02%)
$\blacksquare$ road(14.98%)
$\blacksquare$ parking(2.31%)
$\blacksquare$ sidewalk(6.43%)
$\blacksquare$ other-grnd.(2.05%)
$\blacksquare$ building(15.67%)
$\blacksquare$ fence(0.96%)
$\blacksquare$ vegetation(41.99%)
$\blacksquare$ terrain(7.10%)
$\blacksquare$ pole(0.22%)
$\blacksquare$ traf.-sign(0.06%)
$\blacksquare$ other-struct.(4.33%)
$\blacksquare$ other-obj.(0.28%)
IoU mIoU
Mono-Input Methods
MonoScene (cao2022monoscene) 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
GaussianFormer2 (huang2024probabilistic) 38.31 13.90 21.08 2.55 4.21 12.41 5.73 1.59 54.12 11.04 32.31 3.34 32.01 4.98 28.94 17.33 3.57 5.48 5.88 3.54
TPVFormer (huang2023tri) 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer (zhang2023occformer) 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
IAMSSC (xiao2024instance) 41.80 12.97 18.53 2.45 1.76 5.12 3.92 3.09 47.55 10.56 28.35 4.12 31.53 6.28 29.17 15.24 8.29 7.01 6.35 4.19
Temporal-Input Methods
VoxFormer (li2023voxformer) 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
DepthSSC (yao2023depthssc) 40.85 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.77 5.24 3.51
Symphonies (jiang2024symphonize) 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer (yu2024context) 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
SGN (mei2024camera) 47.06 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86
HD2-SSC (Ours) 48.58 20.62 33.00 0.00 2.98 26.90 12.38 4.75 60.17 17.58 37.55 8.7 40.72 9.51 41.69 14.69 18.7 12.35 16.65 12.75
Table 2: Camera-based 3D semantic scene completion results on the SSCBench-KITTI-360 (behley2019semantickitti) test set. denotes the methods employing EfficientNet-B7 (tan2019efficientnet) as image backbone, and represents the methods utilizing ResNet-50 (he2016deep) as image backbone. Best results are highlights in bold, and second-best scores are underlined.
Experiments
In accordance with existing camera-based SSC methods (mei2024camera; li2023voxformer), extensive experiments and analyses are conducted to validate our HD2-SSC framework on the popular SemanticKITTI (behley2019semantickitti) and SSCBench-KITTI-360 (li2023sscbench) datasets.
Experimental Setup
Dataset.
(1) The SemanticKITTI (behley2019semantickitti) dataset, based on the KITTI Odometry Benchmark (geiger2012we), contains 22 autonomous driving sequences with 20 semantic classes and annotations. We follow the official protocol to split the total sequences into (00-07, 09-10) / (08) / (11-21) for the training / validation / test sets respectively. (2) The SSCBench-KITTI-360 (li2023sscbench) dataset consists of 9 densely annotated sequences of urban driving scenes. It is separated into a training set from sequences (00, 02-05, 07, 10), a validation set from sequence (06), and a test set from sequence (09). For both datasets, the semantic labels are within the range of $0\sim 51.2m,-25.6\sim 25.6m,-2\sim 4.4m]$, and the target voxel grids are represented as $256\times 256\times 32$ voxel grids with the resolution of 0.2mm.
Evaluation Metrics.
Following (mei2024camera), we adopt the Intersection over Union (IoU) of occupied voxels as the evaluation metric for the class-agnostic scene completion (SC) task, and the mean Intersection over Union (mIoU) metric for the semantic scene completion (SSC) task.
Variants IoU mIoU
Baseline 44.15 13.35
Baseline + HSD 46.45 15.58
Baseline + HOR 46.07 16.12
Baseline + HSD + HOR (Ours) 47.59 17.44
Table 3: Ablation study on the SemanticKITTI validation set dataset of different components of our HD2-SSC.
Variants IoU mIoU
HD2-SSC (Ours) 47.59 17.44
w/o $L_{\rm orth}$ 46.93 (-0.66) 16.64 (-0.8)
w/o $L_{\rm decouple}$ 46.85 (-0.74) 16.78 (-0.66)
w/o $L_{\rm critical}$ 46.49 (-1.1) 16.31 (-1.13)
Table 4: Ablation study on the SemanticKITTI validation set dataset of different loss terms.
Implementation Details.
We crop the input RGB images of cam2 to size $1220\times 370$ for SemanticKITTI and images of cam1 to size $1408\times 376$ for SSCBench-KITTI-360, extracting 2D feature maps with $1/16$ of input resolutions for subsequent process. The feature dimension $C$ is set to 128. The size $H\times W\times Z$ of the 3D feature volume is $128\times 128\times 16$, and the final predictions are upsampled to $256\times 256\times 32$. The channel dimensions for 2D pixel features and 3D voxel features are set to $C_{\rm 2D}=256$ and $C_{\rm 3D}=32$, respectively. The expanded dimension in the pseudo voxelization block is $D_{\rm exp}=4$, and the number of pixel and voxel query vectors is $N_{\rm query}=100$. The number of geometric and semantic critical voxels is set as $k=4096$. The above hyperparameters are practically selected by tuning on the validation set. HD2-SSC is trained for 24 epochs on 4 A6000 GPUs with a total batchsize of 4. The AdamW (loshchilov2017decoupled) optimizer is utilized with an initial learning rate of 2e-4 and a weight decay of 1e-2.
Figure 4: Effect of the expanded dimension on the SSC performance, evaluated on the SemanticKITTI validation set.
Figure 5: Visualization results of SSC prediction on the SemanticKITTI validation set. We highlight the occupancy ground truth with blue boxes, false SSC predictions of the best comparison method SGN with red boxes, and the improved SSC predictions from our HD2-SSC approach with green boxes. Better viewed when zoomed in.
Main Results
Quantitative comparison.
[…]
Ablation on architectural components.
To further assess the effectiveness of HD2-SSC, we conduct an ablation study on the SemanticKITTI validation set, evaluating the impact of key components. We incrementally integrate the HSD and HOR modules into the baseline method (li2023voxformer), with results summarized in Table 3. Incorporating HSD yields improvements of 2.30% in IoU and 2.23% in mIoU over the baseline. Adding the HOR module enhances SSC performance over the baseline by 1.92% in IoU and 2.77% in mIoU. These results underscore the benefits of decoupling coarse pixel semantics and refining contextual details with the aligned geometric and semantic distribution. Notably, HSD and HOR modules complement each other, with their combined application achieving superior results.
Ablation on different losses.
Table 4 presents the ablation results on different losses, where disabling each loss causes performance degradation, validating their effectiveness:
  • $L_{\rm orth}$ improves the semantic diversity of expanded 2D features, decoupling coarse pixel semantics.
  • $L_{\rm decouple}$ promotes semantic distinction in semantic clusters, enhancing the aggregated features.
  • $L_{\rm critical}$ aligns critical voxel distributions, improving semantic and geometry consistency in the SSC prediction.
Effect of expanded dimension.
Figure 4 demonstrates the effect of expanded dimension, where the best performance is achieved with $D_{\rm exp}=4$. When the expanded dimension keeps increasing, SSC performance begins to degrade. This phenomenon arises because, without explicit pixel semantic labels, too many expanded dimensions may bring additional confusion of “imaginary” semantics that do not correspond to real-world objects, leading to inferior performance.
Qualitative results.
Figure 5 provides the visualization results from the SemanticKITTI validation set. To better illustrate the effectiveness of our HD2-SSC approach, we also present the visualization results of the best comparison method SGN (mei2024camera) and the corresponding ground truth illustrations. We outline the occupancy ground truth with blue boxes in the second row as a reference for comparison. The red boxes in the fourth row indicate the area of false voxel predictions where the SGN struggles, while the green boxes in the third row highlight the improved scene completion results from our HD2-SSC approach with completion of missing voxels and correction of erroneous ones. This improvement is attributed to HD2-SSC addressing the dimension and density gap with decoupled pixel semantics and refined voxel predictions.
Conclusion
In this paper, we identify the challenges of input-output dimension gap and annotation-reality density gap in camera-based SSC, and introduce the HD2-SSC approach to address them. The HSD module expands and decouples coarse pixel semantics with orthogonal and decoupling loss, mitigating the dimension gap with the aggregation of voxelized semantics. The HOR module aligns geometric and semantic critical voxels in a “detect-and-refine” architecture, promoting consistent contextual details to bridge the density gap. Experiments and analyses on the SemanticKITTI and SSCBench-KITTI-360 datasets validate the effectiveness of our HD2-SSC approach with superior SSC performance.
Limitation. Despite our effectiveness, failure cases such as false prediction and incomplete boundary still occur in severe occlusion and distant areas with confusing features. Our future work aims to incorporate physical regularities to complement low-quality semantic features around those areas.
Acknowledgments
This work was supported by the grants from the National Natural Science Foundation of China (62525201, 62132001, 62432001) and Beijing Natural Science Foundation (L247006, L257005).
Appendix A Appendix
Efficiency Analysis
Table 5 below presents the efficiency evaluation results of our HD2-SSC approach against the best comparison method SGN (mei2024camera). It can be observed that HD2-SSC achieves superior performance with slightly increased model parameters and GFLOPs, even less GPU memory and inference time, since our enhanced feature enables accurate SSC prediction with (128,128,16) feature grids, avoiding SGN’s costly up-sampling to (256,256,32) grids.
Table 5: Efficiency evaluation against the best comparison method SGN on 1 NVIDIA A6000 GPU.
Methods Params (M) Memory (G) GFLOPs Inference (s)
SGN 28.16 15.83 725.05 0.61
HD2-SSC 28.96 14.42 842.76 0.56
Dataset Generalization
Table 6 below presents the evaluation results on the Occ3D-nuScenes (tian2024occ3d) dataset, where our HD2-SSC outperforms popular OccFormer (zhang2023occformer) and BEVDet4D (huang2022bevdet4d) baselines, demonstrating our generalizability across different datasets.
Table 6: Evaluation on the Occ3D-nuScenes dataset.
Methods IoU mIoU
OccFormer 70.1 37.4
BEVDet4D 73.8 39.3
HD2-SSC (Ours) 75.4 44.2
Evaluation with EfficientNet-B7
Table 7 presents the evaluation results with EfficientNet-B7 (tan2019efficientnet) backbone on the SSCBench-KITTI-360 (li2023sscbench) test set, where our HD2-SSC achieves superior performance compared to CGForemr (yu2024context) and L2COcc (wang2025l2cocc), further demonstrating our effectiveness in bridging the dimension and density gaps in the SSC task.
Table 7: Evaluation with EfficientNet-B7 backbone on the SSCBench-KITTI-360 test set.
Methods IoU mIoU
CGFormer 48.07 20.05
L2COcc 48.07 20.11
HD2-SSC (Ours) 49.03 21.13
Failure Case Analysis
Figure 6 presents the failure case analysis on the SemanticKITTI validation set, where failure cases such as false occupancy prediction and incomplete object boundary occur in severe occlusion and distant areas due to the low-quality features with confusing semantics. Our future work aims to incorporate physical regularities to complement low-quality semantic features around those areas.
Figure 6: Failure case analysis on the SemanticKITTI validation set. We highlight the occupancy ground truth with green boxes, false SSC predictions of our HD2-SSC approach with red boxes. Better viewed when zoomed in.